# Learning Semantic Traversability with Egocentric Video and Automated Annotation Strategy

### v1

@@ I Introduction
When people navigate in an urban environment, they usually start by utilizing readily available map services on their smartphones, such as “Google Maps”, to find a rough global route from their current location to their desired destination. They then repeat the process of planning and tracking a local path within their field of view that is most similar to the global path. To ensure safe and responsive closed-loop navigation, not only the similarity with the global path is taken into account when determining the local path, but also the geometric and semantic traversability. Geometric traversability is evaluated using geometric information from the current viewpoint to avoid collisions and rough, bumpy terrains. Semantic traversability is determined by semantic understanding and prior knowledge of the scene, which results in a preference for sidewalks over roads and the use of crosswalks when crossing the street. Humans’ navigation procedures and reasoning capabilities serve as valuable insights for developing an autonomous navigation framework for mobile robots.
@@
Robot navigation frameworks are often built hierarchically with multiple modules, including a mapper, path planner, and controller. The mapper creates a map of the environment in a suitable representation using exteroceptive sensors such as lidar and camera. The path planner then finds a path on the map that maximizes the objective function defined by the engineer. The controller then follows the planned path by executing appropriate high-level (e.g., velocity commands, waypoints)[8] and low-level commands (e.g., joint PD targets)[9]. Similar to how humans navigate, a global route connecting the current location to the desired destination can be obtained from an openly accessible world map. Subsequently, the local path planner adjusts the computed path based on geometric and semantic traversability, which are evaluated and accumulated in the built map[10] during navigation. Geometric traversability is computed using either analytic techniques[11] or learning-based methods[8, 12] from a map constructed with depth cameras or lidar, which contains geometric information such as grid occupancy and terrain height. Semantic traversability is often evaluated using semantic segmentation predictions from semantic segmentation models[13, 14] trained on large-scale datasets containing RGB images and semantic class annotations. Engineers assign a heuristic objective value to each terrain class based on their preference[1, 15, 16]. For example, if the engineer wants the robot to have a preference $Sidewalk>Grass>Road$, the objective values for each terrain class can be set as $Sidewalk=1,\,Grass=0.5,\,Road=0.25$, indicating high values for high preference.
@@
Because incorrect classifications of terrain classes can deteriorate navigation behavior, the trained semantic segmentation models are usually fine-tuned before being deployed with manually collected and annotated data from the testing domain[16]. Concretely, the data is collected by rolling out the robot in the testing domain, often controlled by human experts, and the data annotation is held manually by humans. However, the process of manually preparing the semantic segmentation dataset is highly expensive, labor-intensive, and not scalable. For rapid development of the semantic traversability estimator, a more scalable data collection method and an automated labeling technique are necessary. Furthermore, the model design of the semantic traversability estimator should be adaptable enough to handle OOD terrain classes and OOD visual features.
To this end, we present an effective methodology for training a neural network that predicts semantic traversability using a highly scalable data acquisition method and an automated annotation strategy without the need for manual labeling (Figure 1). Instead of using the target robot to gather data, we utilize egocentric videos obtained by mounting a camera on a person’s chest and capturing the egocentric perspective of a pedestrian crossing an urban area. The collected egocentric videos are then automatically labeled using the Segment Anything Model (SAM)[18], a recent foundation model in image segmentation possessing high generalizability and powerful prompting techniques. The semantically traversable regions in the RGB image are annotated by prompting SAM with approximate footsteps extracted from the egocentric videos using the Monocular Visual-Inertial SLAM[19], eliminating the need for human intervention during the annotation and prompting process. The semantic traversability estimator is then obtained by fine-tuning a lightweight semantic segmentation model[13] with the auto-labeled dataset and minimal model architecture modification. We have conducted extensive quantitative and qualitative evaluations to demonstrate the effectiveness of our method in terms of data collection, data annotation, and model training. Specifically, we show that leveraging egocentric video with the proposed automated annotation strategy possesses great potential for highly scalable and worldwide training data collection due to its minimal hardware requirements. Additionally, we demonstrate that fine-tuning the pre-trained semantic segmentation model yields a semantic traversability estimator that is highly accurate, able to handle diverse camera viewpoints, and lightweight (71Hz on a desktop GPU and 16Hz on an onboard embedded GPU). The trained semantic traversability estimator is further deployed for autonomous quadruped robot navigation in an urban environment, showcasing the method’s capability for real-world applications.
@@ II Related work
There has been various prior research on learning a geometric traversability estimator (typically parameterized with neural networks) in a self-supervised manner without the need for manual labels. The required raw data, such as RGB-D images, was first collected by rolling out the robot in the real-world testing area or the physics simulation[8, 12]. Then, traversability labels were obtained from the human expert’s velocity command[20], velocity command tracking performance[21], collision checking results[8, 12], or IMU sensor data[3, 5]. However, the methods are limited to geometric traversability and place less emphasis on category-level semantic information. Thus, the trained estimators cannot distinguish between terrains that are geometrically similar but semantically different, such as a sidewalk and a flat road.
@@
Several methods have previously been proposed for training a semantic traversability estimator via self-supervised learning. Wellhausen et al.[2] used the future robot footsteps projected on the current image as traversable point labels for training the traversability estimator. Each point label included ground reaction scores from force-torque sensors as well as corresponding terrain classes. Schmid et al.[22] approached the problem as an anomaly detection problem, training an autoencoder to reconstruct only the traversed area of the image. The traversable region on the image was then identified as an area with a low reconstruction error. Frey and Mattamala et al.[4] integrated weak image segmentation masks with automatically derived traversability values from velocity command tracking errors, allowing the traversability estimator to consider semantic cues. There were methods focused on representation learning[23, 24] in which terrain feature encoders were first trained using unsupervised learning and then a binary terrain classifier was trained to determine whether the terrain is preferred or not based on the pre-trained feature. However, these methods necessitate a two-step training process and rely on manual data labeling or heuristics to generate binary terrain preference labels.
Our work is closely related to the ones presented by Wellhausen et al.[2] and Frey and Mattamala et al.[4] but differs from theirs in three ways. First, we obtained traversable area labels from each image by using the recent foundation model in image segmentation (i.e., SAM) and the model’s powerful prompting capability. The proposed annotation method results in large area labels, which are based on visual cues like colors and edges, instead of local point or segment labels. There is a concurrent work done by Jung et al.[25] that used SAM with contrastive learning in self-supervised traversability learning for navigation in field environments. Our work distinguishes itself by its emphasis on urban settings and the integration of semantic information in the annotation process. Second, we utilized egocentric videos taken with a camera mounted on a pedestrian’s chest instead of calibrated sensor modules mounted on the target robot. This minimal hardware requirement enables effective and scalable data collection. Lastly, we tested our approach in an urban setting, which presents a greater challenge for reliable traversability estimation compared to park-like settings or field environments because of the high visual diversity and significance of semantic information.
@@ III-A Data collection
We collect egocentric videos obtained by mounting a GoPro camera on a person’s (i.e., volunteer) chest and capturing the egocentric perspective of the volunteer walking around an urban environment. This approach enables effective and scalable data collection compared to previous works[2, 4, 22, 24, 23] that rely on data obtained directly from the target robot. Leveraging the target robot itself often requires well-calibrated sensor settings and limits the data acquisition capability based on several constraints such as the number of available robots, the robot’s speed, and physically reachable area. The volunteer is instructed to walk around inside the designated area while taking into account their visual path preference, which is consistent with urban regulations such as pedestrians walking on the sidewalk rather than the road and using crosswalks when crossing the road. The volunteer is not restricted from going to specific terrains, such as steep bumps, because our target robot for deploying the semantic traversability estimator is a legged robot that can traverse most regions that humans can traverse. However, volunteers may be given instructions in advance to avoid certain terrains, taking into account the type of target robot and its controller capability, which prevents the corresponding terrain from being labeled as traversable. For example, if the target robot is a wheeled robot, the volunteer may be told to avoid stairs and vertical slopes because the robot cannot overcome them.
The methodology of leveraging egocentric videos obtained from humans for robot navigation is inspired by the Ego4D dataset[26], a massive-scale egocentric video dataset capturing diverse daily-life activities. The work[26] demonstrated the possibility of collecting egocentric videos on a large scale, and the following works showed the usage of egocentric videos in robotic tasks such as affordance learning for robot manipulation[27]. Our work aims to take advantage of egocentric videos in robot navigation and use them to learn semantic traversability.
@@ III-B Automatic annotation
We use a large semantic segmentation model [14] in conjunction with SAM to produce the final semantic traversability labels. This process introduces more detailed semantic information required for urban environment deployment (e.g., road is less preferred compared to other traversable terrains) and can be removed if not necessary (e.g., park-like environment). Even though a large semantic segmentation model’s computational load makes it unsuitable for high-frequency real-time usage (i.e., Mask2former[14] model that we used runs at 1.66Hz on the onboard embedded GPU), this is less important for offline data annotation. From the semantic segmentation predictions, segments labeled as crosswalk and road are added to and removed from the SAM mask predictions, respectively. If projected footsteps are sufficiently included in the remaining masks, the remaining segmented areas are annotated with a value of 1. If not, this is typically a scenario when the volunteer walked on the road because it was the best navigation behavior in their current field of view. In this instance, the areas where the SAM mask intersects with areas identified as road are given a value of 0.25, whereas the remaining SAM mask areas and areas identified as crosswalk are given a value of 1. The annotation process is summarized in Figure 2.
@@ IV-B Evaluation of the trained semantic traversability estimator
TABLE I: Quantitative evaluation results of semantic traversability estimators
@@
We also compared the performance with variations of our method. Specifically, the model backbone was trained from scratch, the model backbone was frozen and only the adaptor was trained, the adaptor architecture was converted to a two-layer U-Net[28], and the model backbone was converted to MobileSAM[29], a compressed version of SAM through knowledge distillation. MobileSAM was chosen as an alternative model backbone over SAM because SAM requires more VRAM to fine-tune than the ones we have (NVIDIA GeForce RTX 4090 (24GB)) and is not suitable for real-time inference on an onboard embedded GPU due to its high computational cost.
@@
Figure 5: Estimation results of semantic traversability estimators
@@
Table I and Figure 5 present the quantitative and qualitative evaluation results, respectively. Our traversability estimator, which was obtained by fine-tuning a lightweight semantic segmentation model with area labels, outperformed other methods across all metrics. It demonstrated the highest accuracy in determining traversable areas, taking into account the highest precision, recall, and IoU values, as well as the lowest RMSE. Furthermore, it demonstrated the shortest computation time, allowing for semantic traversability estimation at a high frequency of 71Hz on a desktop GPU and 16Hz on an onboard embedded GPU. The model trained with only point labels, as in [2], showed low IoU values and only local area predictions due to small supervision signals during training, even with strong data augmentation. This demonstrates the significance of our automatic annotation method, which expands point labels to area labels through SAM, for more efficient training of image data with high visual diversity. Reconstruction-based methods [22] showed relatively low performance with large incorrect predictions due to the it’s limited capability to reconstruct high-frequency information (e.g., sidewalk patterns) and discriminate traversable regions based on reconstruction errors. Directly obtaining traversability values from semantic segmentation models showed lower accuracy due to incorrect predictions for non-dominant categories in the large-scale dataset the models were originally trained on, as well as high computation time due to a large number of neural network parameters and the computationally expensive argmax operation required for category-wise value assignments.
@@
We also demonstrate the effectiveness of our training methodology, which involves leveraging pre-trained semantic segmentation model weights and fine-tuning them end-to-end. As shown in Table I, models that do not use pre-trained weights or only fine-tune the added layers showed lower accuracy regardless of the added layers’ size and neural network architecture. Interestingly, the model initialized with a segmentation model (i.e., MobileSAM [29]) showed lower accuracy than ours initialized with a semantic segmentation model (i.e., Segformer [13]). This implies that pre-trained representations of the semantic segmentation model are more suitable for semantic traversability estimation than those in the segmentation model and that fine-tuning them end-to-end is critical for adapting the OOD RGB features and terrain classes.

### v2

@@ I Introduction
Robot navigation frameworks are often built with several modules, one of which is the path planner. The path planner finds a path on the map[8] by considering the geometric and semantic traversability of each point. Geometric traversability is computed using either analytic techniques[9] or learning-based methods[10, 11] from a map constructed with depth cameras or lidar, which contains geometric information such as grid occupancy and terrain height. Semantic traversability is often evaluated using semantic segmentation predictions from semantic segmentation models[12, 13] trained on large-scale datasets containing RGB images and semantic class annotations. Engineers assign a heuristic objective value to each terrain class based on their preference[1, 14, 15]. For example, if the engineer wants the robot to have a preference $Sidewalk>Grass>Road$, the objective values for each terrain class can be set as $Sidewalk=1,\,Grass=0.5,\,Road=0.25$, indicating high values for high preference.
@@
Because incorrect classifications of terrain classes can deteriorate navigation behavior, the trained semantic segmentation models are usually fine-tuned before being deployed with manually collected and annotated data from the testing domain[15]. However, the process of manually preparing the semantic segmentation dataset is highly expensive, labor-intensive, and not scalable. For rapid development of the semantic traversability estimator, a more scalable data collection method and an automated labeling technique are necessary. Furthermore, the model design of the semantic traversability estimator should be adaptable enough to handle OOD terrain classes and OOD visual features.
To this end, we present an effective methodology for training a neural network that predicts semantic traversability using a highly scalable data acquisition method and an automated annotation strategy without the need for manual labeling (Figure 1). Instead of using the target robot to gather data, we utilize egocentric videos obtained by mounting a camera on a person’s chest and capturing the egocentric perspective of a pedestrian crossing an urban area. The collected egocentric videos are then automatically labeled by prompting a segmentation model[17] with approximate footsteps extracted using the Monocular Visual-Inertial SLAM[18] and refining the outputs with a large semantic segmentation model[13]. The semantic traversability estimator is then obtained by fine-tuning a lightweight semantic segmentation model[12] with the auto-labeled dataset and minimal model architecture modification. We have conducted extensive quantitative and qualitative evaluations to demonstrate the effectiveness of our method in terms of data collection, data annotation, and model training. Specifically, we show that leveraging egocentric video with the proposed automated annotation strategy possesses great potential for highly scalable and worldwide training data collection due to its minimal hardware requirements. Additionally, we demonstrate that fine-tuning the pre-trained semantic segmentation model yields a semantic traversability estimator that is highly accurate, able to handle diverse camera viewpoints, and lightweight (71Hz on a desktop GPU and 16Hz on an onboard embedded GPU). The trained semantic traversability estimator is further deployed for autonomous quadruped robot navigation in an urban environment, showcasing the method’s capability for real-world applications.
@@ II Related work
There have been several prior research on learning a semantic traversability estimator (typically parameterized with neural networks) in a self-supervised manner without the need for manual labels. Wellhausen et al.[2] used the future robot footsteps projected on the current image as traversable point labels for training the traversability estimator. Each point label included ground reaction scores from force-torque sensors as well as corresponding terrain classes. Schmid et al.[19] approached the problem as an anomaly detection problem, training an autoencoder to reconstruct only the traversed area of the image. The traversable region on the image was then identified as an area with a low reconstruction error. Frey and Mattamala et al.[4] integrated weak image segmentation masks with automatically derived traversability values from velocity command tracking errors, allowing the traversability estimator to consider semantic cues. Jung et al.[20] used Segment Anything model (i.e., SAM)[17] and contrastive learning in self-supervised traversability learning for field environment navigation.
Our work is closely related to several previous approaches but greatly distinguishes itself in two points. First, we leverage a two-step annotation method in which the labels are first produced from the point-prompt-based segmentation model (i.e., SAM) and then refined with the large semantic segmentation model (i.e., Mask2former). The process results in large traversable area labels instead of local point[2] or segment[4] labels and enables larger supervision signal during training. Furthermore, the additional refinement step allows for more fine-grained and reliable traversability labels by incorporating semantic information and filtering out SAM’s overestimated predictions[20], which are frequently encountered due to a lack of visual cues. Our extensive evaluations under a variety of environmental conditions demonstrate the effectiveness of the proposed two-step annotation method, as well as its applicability for use in not only field and park-like environments but also urban scenarios. Second, we utilize egocentric videos captured by pedestrians as the data source and demonstrate its capability for the usage of training mobile robots’ traversability estimators. The suggested data acquisition method enables effective and scalable data collection compared to prior research [2, 4, 20, 19], which collected data directly by deploying the target robot. To the best of our knowledge, our work is the first to demonstrate that high-performing traversability estimators can be trained with egocentric videos recorded by pedestrians and further deployed for real-robot navigation in urban scenarios.
@@ III-A Data collection
We collect egocentric videos obtained by mounting a GoPro camera on a person’s (i.e., volunteer) chest and capturing the egocentric perspective of the volunteer walking around an urban environment. This approach enables effective and scalable data collection compared to previous works[2, 4, 19, 21, 22] that rely on data obtained directly from the target robot. Leveraging the target robot itself often requires well-calibrated sensor settings and limits the data acquisition capability based on several constraints such as the number of available robots, the robot’s speed, and physically reachable area. The volunteer is instructed to walk around inside the designated area while taking into account their visual path preference, which is consistent with urban regulations such as pedestrians walking on the sidewalk rather than the road and using crosswalks when crossing the road. The volunteer is not restricted from going to specific terrains, such as steep bumps, because our target robot for deploying the semantic traversability estimator is a legged robot that can traverse most regions that humans can traverse[23]. However, volunteers may be given instructions in advance to avoid certain terrains, taking into account the type of target robot and its controller capability, which prevents the corresponding terrain from being labeled as traversable. For example, if the target robot is a wheeled robot, the volunteer may be told to avoid stairs and vertical slopes because the robot cannot overcome them.
@@ III-B Automatic annotation
We use a large semantic segmentation model [13] in conjunction with SAM to produce the semantic traversability labels for the training dataset. This process introduces more detailed semantic information required for urban environment deployment (e.g., road is less preferred compared to other traversable terrains) and filters out SAM’s overestimated predictions. Even though a large semantic segmentation model’s computational load makes it unsuitable for high-frequency real-time usage (i.e., Mask2former[13] model that we used runs at 1.66Hz on the onboard embedded GPU), this is less important for offline data annotation. From the predictions of Mask2former, segments labeled as crosswalk and road are added to and removed from the SAM mask predictions, respectively. If projected footsteps are sufficiently included in the remaining masks, the remaining segmented areas are annotated with a value of 1. If not, this is typically a scenario when the volunteer walked on the road because it was the best navigation behavior in their current field of view. In this instance, the areas where the SAM mask intersects with areas identified as road are given a value of 0.25, whereas the remaining SAM mask areas and areas identified as crosswalk are given a value of 1. The annotation process is summarized in Figure 2.
@@ IV-B Evaluation of the trained semantic traversability estimator
TABLE I: Quantitative evaluation results on the urban dataset
@@
We also compared the performance with variations of our method. Specifically, the model was fine-tuned with labels acquired without the refinement stage (i.e., only SAM is used for annotation), the model backbone was trained from scratch, the model backbone was frozen and only the adaptor was trained, the adaptor architecture was converted to a two-layer U-Net[24], and the model backbone was converted to MobileSAM[25], a compressed version of SAM through knowledge distillation. MobileSAM was chosen as an alternative model backbone over SAM due to its low computational cost.
Figure 5: Estimation results of semantic traversability estimators
@@
Table I and Figure 5 present the quantitative and qualitative evaluation results, respectively. Our traversability estimator, which was obtained by fine-tuning a lightweight semantic segmentation model with area labels, outperformed other methods across all metrics. It demonstrated the highest accuracy in determining traversable areas, taking into account the highest precision, recall, and IoU values, as well as the lowest RMSE. Furthermore, it demonstrated the shortest computation time, allowing for semantic traversability estimation at a high frequency of 71Hz on a desktop GPU and 16Hz on an onboard embedded GPU. The model trained with only point labels, as in [2], showed low IoU values and only local area predictions due to small supervision signals during training. Concretely, the training signal is limited to dealing with extremely visually diverse image data collected in urban environments because the area indicated as traversable is significantly smaller than the area labeled as non-traversable. This demonstrates the significance of our automatic annotation method, which expands point labels to area labels through SAM, for more efficient training of image data with high visual diversity. Reconstruction-based methods [19] showed relatively low performance with large incorrect predictions due to the it’s limited capability to reconstruct high-frequency information (e.g., sidewalk patterns) and discriminate traversable regions based on reconstruction errors. Directly obtaining traversability values from semantic segmentation models showed lower accuracy due to incorrect predictions for non-dominant categories in the large-scale dataset the models were originally trained on, as well as high computation time due to a large number of neural network parameters and the computationally expensive argmax operation required for category-wise value assignments.
The traversability estimator trained using labels obtained without the refinement step performed poorly compared to those obtained with it. This is because the refinement step improves the quality of labels from SAM by converting the binary labels to fine-grained labels and removing inflated parts based on semantic information (Figure 4, Figure 8).
@@
We also demonstrate the effectiveness of our training methodology, which involves leveraging pre-trained semantic segmentation model weights and fine-tuning them end-to-end. As shown in Table I, models that do not use pre-trained weights or only fine-tune the added layers showed lower accuracy regardless of the added layers’ size and neural network architecture. Interestingly, the model initialized with a segmentation model (i.e., MobileSAM [25]) showed lower accuracy than ours initialized with a semantic segmentation model (i.e., Segformer [12]). This implies that pre-trained representations of the semantic segmentation model are more suitable for semantic traversability estimation than those in the segmentation model. In our experience, representations of segmentation models (e.g., SAM, MobileSAM) are more focused on visual cues such as colors or edges rather than semantic information. Our experimental results indicate that fine-tuning the semantic segmentation model end-to-end is critical for adapting the OOD RGB features and terrain classes. Furthermore, leveraging the semantic segmentation model’s pre-trained model weights allows for effective training with a small quantity of data and acts as a weak inductive bias for traversability estimation.
@@ IV-C Evaluation of real-robot navigation
IV-D Evaluation of open-source dataset
To show the applicability of our method across different domains from our custom dataset, we conducted an analysis using the Rellis3D dataset[26]. The Rellis3D dataset is an open-source dataset constructed by rolling out a wheel-based mobile robot in field environments. To apply our method to the Rellis3D setting, the time horizon $T^{\prime}$ is modified from 3s to 4.5s, and segments labeled as “vegetation” are removed from the SAM mask predictions rather than “road” or “crosswalk”. These changes are made due to the difference in vehicle speed and application domain from urban to field settings. Other engineering choices and hyperparameters for both the automatic annotation and training are kept identical to those utilized for our custom dataset. For evaluation, we employed the same process and criteria as for our own dataset. Category-wise heuristic traversability values for ”Segformer + Heuristic”333[Path, Grass, Sand, Dirt track, Land, Field]: 1, [Hill, Earth]: 0.5 and ”Mask2former + Heuristic”444[Terrain]: 1 were updated to reflect the different environment settings.
TABLE II: Quantitative evaluation results on the rellis3d dataset
Precision Recall IoU RMSE
[Segformer + 1x1 conv]
+ Area labels (Ours)
0.922
(0.853 / 0.990)
0.953
(0.977 / 0.929)
0.834 0.212
[Segformer + 1x1 conv]
+ Point labels [2]
0.867
(0.975 / 0.758)
0.657
(0.316 / 0.997)
0.312 0.460
Segformer + Heuristic [1, 14, 15]
0.819
(0.652 / 0.985)
0.870
(0.967 / 0.773)
0.640 0.395
Mask2former + Heuristic [1, 14, 15]
0.919
(0.868 / 0.971)
0.935
(0.930 / 0.939)
0.807 0.232
[Segformer + 1x1 conv]
+ Area labels w/o refinement
0.871
(0.748 / 0.994)
0.912
(0.986 / 0.837)
0.737 0.308
Figure 8: Qualitative results on Rellis3D dataset. The top two lines display the stepwise annotation results, while the bottom two lines show the estimated outputs from the trained traversability estimator. In Mask2former predictions, only terrains and vegetations are color-segmented with brown and purple if detected.
The quantitative and qualitative evaluation results on the Rellis3D dataset are shown in Table II and Figure 8. Our traversability estimator performed better than other approaches on all metrics, which is consistent with the experimental results exhibited in our custom dataset. Furthermore, our two-step automatic annotation pipeline enables to generate clear traversability labels on areas where either SAM or Mask2former alone cannot accurately predict. Interestingly, the introduction of the semantic refinement stage in the annotation process significantly improved the traversability estimator’s performance when compared to urban situations. This is because, in field situations, SAM frequently generates inflated labels due to ambiguous visual cues. The filtering process based on the predictions from the large semantic segmentation models enables to alleviate the overestimation problem of SAM and produce more reliable labels.
Our experimental results and analysis on the Rellis3D dataset show that our methodology is not limited to urban settings or our custom dataset, but can also be easily extended to other environment settings (e.g., field, park) or open-source datasets, allowing for the training of highly accurate and computationally light semantic traversability estimators.
@@ V Conclusion
Figure 9: Limitations of the proposed semantic traversability estimator. (Left) Occasionally, speed bumps are misidentified as crosswalks, leading to incorrect predictions of traversability. (Right) Crosswalks located at a distance are sometimes not recognized as traversable.
@@
As shown in Figure 9, the proposed semantic traversability estimator sometimes encounters limitations when visual features are not clearly identifiable, such as with similarly appearing objects or distant objects. However, these limitations do not significantly impact navigation scenarios, as the visual features become clearer as the robot approaches.